OpenStudy (anonymous):

xsquared-2x-13=0 ising the india method

OpenStudy (shayaan_mustafa):

Hi lindy56 :) Here is you requirement dear. A. Move the constant term to the right side of the equation. B. Multiply each term in the equation by four times the coefficient of the x^2 term. C. Square the coefficient of the original x term and add it to both sides of the equation. D. Take the square root of both sides. E. Set the left side of the equation equal to the positive square root of the number on the right side of the equation and solve for x. Good luck. :)

OpenStudy (anonymous):

i have tried that but i get so far and i get confused to what am doing

OpenStudy (shayaan_mustafa):

At what step you stuck?

OpenStudy (anonymous):

i am not sure if i am even ding it right i xsquared-2x=13 4xsuared -8x=52 4xsquared-8x=56

OpenStudy (anonymous):

i am not sure if i am on the right track

OpenStudy (shayaan_mustafa):

Here you did wrong in the third step. Read step C \[\large 4x^2-8x+4=56+4\] \[\large 4x^2-8x+4=60\] Got it?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thank you

OpenStudy (shayaan_mustafa):

Welcome.